P(AIB) PLA IBC) 21 Pla)-P(B) = PIA OB)

how do i prove this?
Answer #1

Proof [ PLAIB) -1 0 PLAIBO PLAIB) = P(AMB) , P(AIB) =P(AMB) 7 P(3) S PIB) PCOO= 1-PCB) , PeanBo = P(A)-PAANB) 7 P(A1B) = P(A/